OpenStudy (anonymous):

What is the sum of all the elements of the 9th term, which has 9 elements in it, of the following sequence: (1), (3,5), (7,9,11), (13,15,17,19), ⋯?

OpenStudy (amistre64):

well, the terms seem to be 2 apart in each element

OpenStudy (amistre64):

1 3 7 13 2 4 6 im going with add 8 next

OpenStudy (anonymous):

Note that we are basically summing up all the terms from n = 37 to n = 45 in the arithmetic sequence: \(\bf a_n=1+(n-1)2\). But we can define our arithmetic sequence such that it starts at the 37th term. We know that \(\bf a_{37}=73\) so we can define a new arithmetic sequence as \(\bf a_n=73+(n-1)2\) and sum up the first 9 terms of this sequence using the arithmetic sum formula \(\bf S_n=\frac{n}{2}(2a+n-1(d))=\frac{n}{2}(146+(n-1)2)\). Evaluate the sum for n = 9 and that will be your answer.

OpenStudy (anonymous):

@denis1206

OpenStudy (anonymous):

corect